A student is asked to make 184mL of a 0.6 M NaOH solution. How many grams of NaOH will the student need?

1
Expert's answer
2021-05-17T04:20:05-0400

"m(NaOH)=184mL\times\frac{0.6mol(NaOH)}{1000mL}\times\frac{40.0g (NaOH)}{1mol(NaOH)}=4.42g"

Since the least precise measurement has 1 significant digit (0.6 M), the final answer should be reported to 1 significant digit, too. Therefore, it is 4 grams.


Answer: 4 g